OpenStudy (anonymous):

Medals are given! Calculate the volume around the y-axis, limited by y=sqrt(x) and x=4. Thank you.

OpenStudy (anonymous):

I got to \[V= \int\limits_{0}^{4} \pi* (y)^2 dx\] Am I thinking correct?

OpenStudy (anonymous):

= \[\int\limits_{0}^{4} \pi* \left(\begin{matrix}4^3 \\3\end{matrix}\right) = 67.02 ae\]

OpenStudy (kagıtucak):